# SEE-2-SOUND  [Uncaptioned image] : Zero-Shot Spatial Environment-to-Spatial Sound

### v1

@@ Abstract
Figure 1: SEE-2-SOUND enables spatial audio creation from real or generated visual content. Our method takes in an image or a sequence of frames (video) and generates spatial audio with a variable number of different sources. The generated audio could also be conditioned on a text prompt.
1 Introduction
@@
Human visual and auditory systems inherently depend on synchronized spatial audio cues for an immersive experience and contextual understanding of video content. Modern generative models have been able to create compelling high-resolution digital content across various modalities. These models have been successful in generating images [61, 58], text [55, 1], speech [75, 57], audio [44, 35], 3D scenes [40, 53], 4D (dynamic 3D) scenes [68, 42], and videos [67, 13]. Although generating audio from text [30, 18] - and more recently, images [72] - has become a reality, generating spatial audio is still a challenge. Unlike these earlier works, we are interested in solving the specific problem of generating high-quality spatial audio or surround sound to accompany images, animated images (GIFs), and videos. Such a model could be particularly useful to pair with any kind of video generation or image generation techniques as well as on real images and videos.
@@
Audio generation is a long-standing challenge, often subdivided into generating speech, music, and sound effects. Most audio generation methods include inductive biases for each of these tasks, often due to the diversity between the generation of speech, music, and other types of audio [11, 2, 43, 70]. There have also been newer approaches that develop unified representations [44] to generate complex audio as well. However, none of these currently emphasize the generation of rich spatial audio. The few related works that perform image-to-audio [72, 71] are closest to our work. However, these also face the same problem of not being able to generate rich spatial audio. Currently, there also exist a few learned approaches that generate spatial audio [47, 41] but these are often based on having access to the captured audio; following which these approaches localized the audio in a viewing sphere to produce spatial audio for a captured video. However, none of these approaches aim to generate accompanying spatial audio for images, animated images (GIFs), and videos, which remains an open problem.
@@
Towards this goal, we propose SEE-2-SOUND, a training-free method that approaches this specific problem by employing the foundational models, CoDi [72], DepthAnything [84], and Segment Anything [33]. The pipeline works by explicitly factorizing this generative problem into separate problems: (1) identifying regions of interest in the input; (2) finding the 3D location on the viewing sphere for a visual element; (3) generating audio from the regions of interest, and (4) simulating these multiple mono-audio outputs to generate spatial audio. We then explicitly compose these component models to generate spatial audio conditioned on the visual input (e.g. images, animated images (GIFs), or videos) as shown in Figure 1. As our method is compositional and zero-shot, we hope to see SEE-2-SOUNDwidely used or adapted for content creation and enhancement as well as accessibility.
@@
By generating spatial audio that accurately corresponds to visual elements, we can significantly enhance the immersion of images and videos, creating more realistic and engaging user experiences with digital content.SEE-2-SOUND presents a step towards true complete video and image generation, by also incorporating non-verbal audio that accompanies visual generation models [67, 13, 8, 9, 46, 79]. Additionally, our pipeline can also be employed on real visual content.
Our contributions can be summarized as follows:
@@ 2 Related works
The audio generation field has been gaining much attention lately, primarily in models that create audio conditioned on different modalities. From speech synthesis to music generation and sound effects creation, past efforts have explored numerous aspects of auditory information synthesis. Learning algorithms have led to ever-improving techniques in the models for generating high-quality, diverse audio. This surge in interest has fueled research efforts that push the boundaries of audio generation.
2.1 Conditional Audio Generation
@@
Recent work in conditional audio generation has made significant progress using latent diffusion models [43] such as AudioLDM (2) [43, 44] and Make-An-Audio [30]. Audiogen [35] was one of the first to create a text-to-audio environment, treating audio generation as a conditional language modeling task. Building on this, AudioLDM-2 [44] introduces a unified audio generation framework that addresses different objectives for speech, music, and sound effects, using the “language of audio" (LoA) and a GPT-2 [54] model for translation and latent diffusion [58] for generation.
@@
AudioLM [10] employs neural codecs [78] for long-term consistency in audio generation, mapping input audio to discrete tokens and using a language modeling task. Meanwhile, DiffSound [83], using a text encoder, decoder, vector quantized variational autoencoder (VQVAE) [32], and “vocoder”, introduces a mask-based text generation strategy (MBTG) for generating text descriptions from audio labels to address the scarcity of audio-text paired data. NaturalSpeech [70] integrates a memory component in variational autoencoders [32] to achieve human-level quality in text-to-speech systems.
@@
Significant progress in text-to-music generation has been made by notable models such as MusicLM [2], Noise2Music [29], MusicGen [18] and MeLoDy [37]. MusicLM [2] uses contrastive pretraining to align music and language embeddings, improving performance without explicit text using a semantic modeling stage based on w2v-BERT [17]. MusicGen [18] adopts a language modeling approach with melodic conditioning, enhancing control over output. MeLoDy [37] employs dual-path diffusion-based modeling for computational efficiency in music generation.
@@
The image-to-audio and video-to-audio domains have also seen significant advancements. For instance, Im2Wav [66] generates audio from images or image sequences, using Transformers [77] to manipulate hierarchical discrete audio representation derived from VQ-VAE [76]. Similarly, based on GANs [26]: SpecVQGAN [31] enhances visually driven audio synthesis from videos, while RVQGAN [36] achieves high-fidelity audio compression. [7] introduces aligned representations that help in cross-modal retrieval and transferring between audio, visual, and text modalities. Furthermore, AV-NeRF [39] synthesizes new videos with spatial audio along different camera trajectories, and AV-RIR [56] estimates audio-visual room impulse responses from images and audio.
@@
Multi-modal generative models of note include CoDi(-2) [9, 71]. CoDi [9] generates multiple output modalities from various input combinations using a composable generation strategy. CoDi-2 [71] extends this with enhanced modality-language alignment for encoding and generative tasks, improving zero-shot capabilities in multi-modal generation, reasoning, and interactive conversations.
@@
Recent works like MAGNeT [93] introduce a masked generative sequence modeling method for direct audio token generation, achieving higher efficiency in text-to-music and text-to-audio tasks. DIFF-FOLEY [45] improves video-to-audio synthesis quality, relevance, and synchronization using latent diffusion and contrastive audio-visual pretraining. Affusion [82] enhances text-to-audio generation by leveraging text-to-image diffusion models, producing high-quality audio closely aligned with text prompts. These advancements significantly expand the possibilities of audio generation.
2.2 Acoustic Matching
@@
Recent advancements in acoustic matching integrate visual and audio cues to enhance realism and immersion in generated content. AViTAR [14] performs visual-acoustic matching using a cross-modal transformer to integrate visual properties into audio and learning from in-the-wild videos. AdVerb [16] uses a geometry-aware cross-modal transformer to enhance reverberation by leveraging visual cues, significantly outperforming traditional methods. SoundSpaces 2.0 [15] provides geometry-based audio rendering for 3D environments, supporting tasks like audio-visual navigation and source localization with high fidelity and realism.
2.3 Construction of Surround Sound
@@
In 1881, Ader used a pair of microphones and telephone receivers to relay the sound signals at the Paris Opera [60], a first-of-its-kind spatial (stereophonic) sound demonstration. In the past three decades, the development of digital signal processing [34], computer, and internet technologies have rapidly evolved spatial sound [81]. Together, they have made spatial audio commonplace now [88].
@@
In the context of virtual environments generated by game engines, spatial audio is crucial to provide an immersive experience [12]. Traditionally, sounds in virtual worlds are created by recording individual sound sources separately and then blending them according to the intended scene setup [48] using various audio processing techniques [25, 19]. But this recording process is costly [88] and optimizing it involves generating spatial audio. While spatializing mono-audio is a popular area of research (discussed below), to the best of our knowledge, there have not been any approaches to generate spatial audio directly.
2.4 Audio Spatialization
@@
Recent advancements in audio spatialization have focused on integrating audio and visual modalities to enhance realism. Sep-Stereo [91] uses an associative pyramid structure to generate stereo output by combining audio and visual information. Similarly, LAVSS [85] introduces a location-guided audio-visual spatial audio separator, highlighting the importance of visual cues in audio generation.
@@
Several studies have synchronized visual cues and encoders to produce natural sound for dynamic video scenarios [92]. Methods like those in [22, 23] use a U-Net to encode monaural input and decode binaural output with visual guidance.
@@
Other approaches, such as  [21, 35] generate audio samples conditioned on diverse inputs such as text, motion key points, and position information [24, 51], respectively. Techniques like [47] utilize end-to-end trainable neural networks for sound source separation and localization in $360$-degree videos. Additionally, works like [38] synthesize directional impulse responses by exploiting the spatial and directional characteristics of room reverberation.
2.5 Audio Visual Learning
@@
In recent years, there has been a growing interest in audio-visual learning, focusing on understanding the inherent connection between visual and auditory stimuli within vision research. Noteworthy works have delved into explaining the natural synchronization between visual and audio elements [4, 6, 50]. This approach finds application in diverse domains, including speech recognition utilizing both audio and visual cues [69, 27, 86, 90], event localization through joint analysis of audio and visual information [74, 73, 80], and sound source localization [5, 28, 59, 64, 74, 89]. Additionally, there have been explorations in the use of audio-visual data for unsupervised representation learning [49, 50].
@@
3 Preliminary: Conditional LDM
Latent Diffusion models (LDMs) [58] project input images into a lower-dimensional latent space. Unlike traditional diffusion models that operate in a high-dimensional pixel space, LDMs perform forward and backward diffusion processes in a lower-dimensional learned latent space. For generating audio, we have a conditional LDM ${\bm{\epsilon}}_{\theta}$, which can be interpreted as a sequence of denoising autoencoders ${\bm{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\ldots)$; $t\sim\mathcal{U}({1,\ldots,T})$, trained using a squared error loss (representing the denoising term in the ELBO) to denoise some variably noisy audio signals using the following objective,
$\mathbb{E}_{\mathcal{E}(\mathbf{x}),\mathbf{c},{\bm{\epsilon}}\sim\mathcal{N}(0,I),t}\!\left[|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\tau_{\theta}(\mathbf{c}))|^{2}_{2}\right],$ (1)
@@ 3 Preliminary: Conditional LDM
4 Methodology
Figure 2: We show an overview of how our approach works. First, we perform image segmentation and find regions of interest for different objects in the image. We also perform monocular depth estimation on the image and produce 3D co-ordinate estimating positions for each object in the scene. We then generate mono-audio from the regions of interest. Finally, we place all of these audio according to the 3D co-ordinates on a viewing sphere producing spatial or surround sound.
@@ 4.1 Source Estimation
We first focus on finding out regions of interest from the image, $\mathbf{I}$. We do this by finding a segmentation mask ${\mathcal{M}}\in{\mathbb{R}}^{H\times W}$, using a pre-trained Segment Anything model [33]. From the segmentation mask ${\mathcal{M}}$, we extract the contours of each region of interest. Let ${\mathcal{C}}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{n}\}$ represent the set of contours, where each contour $\mathbf{c}_{k}$ is a sequence of points outlining one region of interest. We then compute the smallest polygon ${\mathcal{P}}_{k}$ that bounds each contour $\mathbf{c}_{k}$. Formally, we solve the following optimization problem for each contour,
$\min_{{\mathcal{P}}_{k}}\operatorname{Area}({\mathcal{P}}_{k})\text{ , subjectto, }\forall\mathbf{c}_{k}\in{\mathcal{P}}_{k},$ (2)
@@
${\mathcal{R}}_{k}=\left[\min(\mathbf{c}_{k}^{(x)}),\max(\mathbf{c}_{k}^{(x)}),\min(\mathbf{c}_{k}^{(y)}),\max(\mathbf{c}_{k}^{(y)})\right].$ (3)
@@
For the image $\mathbf{I}$, we also estimate the monocular depth map ${\mathcal{D}}\in{\mathbb{R}}^{H\times W}$ using a pre-trained Depth Anything model [84]. Now we can construct a set of image patches $\mathbf{I}_{\mathbf{p}}$, for the regions of interest from either ${\mathcal{P}}$ or ${\mathcal{R}}$ to create a set of variably sized patches, and the 3D coordinates $(x,y,z)$, from the image coordinates and the monocular depth coordinates illustrated in Figure 2. We find that using AABBs usually works better for generating spatial audio.
4.2 Audio Generation

### v2

# SEE-2-SOUND: Zero-Shot Spatial Environment-to-Spatial Sound

@@ Abstract.
Spatial Audio Generation, Zero-Shot Learning, Generative Models
copyright: rightsretainedjournalyear: 2025conference: SIGGRAPH Posters ’25; August 10-14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Posters (SIGGRAPH Posters ’25), August 10-14, 2025doi: 10.1145/3721250.3742965
@@
Figure 1. SEE-2-SOUND enables spatial audio creation from visual content. Our method takes in an image or a video and generates spatial audio with a variable number of different sources. The audio could also be conditioned on a text prompt.
[…]
Showing an example input to and output from our framework.
1. Introduction
@@
Human visual and auditory systems inherently depend on synchronized spatial audio cues for an immersive experience and contextual understanding of video content. Modern generative models have been able to create compelling high-resolution digital content across various modalities. These models have been successful in generating images (Saharia et al., 2022; Rombach et al., 2022), text (Raffel et al., 2020; Achiam et al., 2023), speech (van den Oord et al., 2016; Ren et al., 2021), audio (Liu et al., 2024; Kreuk et al., 2023), 3D scenes (Lin et al., 2023; Poole et al., 2023), 4D (dynamic 3D) scenes (Singer et al., 2023b; Ling et al., 2023), and videos (Singer et al., 2023a; Brooks et al., 2024). Although generating audio from text (Huang et al., 2023a; Copet et al., 2024) - and more recently, images (Tang et al., 2024) - has become a reality, generating spatial audio is still a challenge. Unlike these earlier works, we are interested in solving the specific problem of generating high-quality spatial audio or surround sound to accompany images, animated images (GIFs), and videos. Such a model could be particularly useful to pair with any kind of video generation or image generation techniques as well as on real images and videos.
@@
Audio generation is a long-standing challenge, often subdivided into generating speech, music, and sound effects. Most audio generation methods include inductive biases for each of these tasks, often due to the diversity between the generation of speech, music, and other types of audio (Bresin et al., 2010; Agostinelli et al., 2023; Liu et al., 2023; Tan et al., 2024). There have also been newer approaches that develop unified representations (Liu et al., 2024) to generate complex audio as well. However, none of these currently emphasize the generation of rich spatial audio. The few related works that perform image-to-audio (Tang et al., 2024, 2023) are closest to our work. However, these also face the same problem of not being able to generate rich spatial audio. Currently, there also exist a few learned approaches that generate spatial audio (Morgado et al., 2018; Lin and Wang, 2021) but these are often based on having access to the captured audio; following which these approaches localized the audio in a viewing sphere to produce spatial audio for a captured video. However, none of these approaches aim to generate accompanying spatial audio for images, animated images (GIFs), and videos, which remains an open problem.
@@
Towards this goal, we propose SEE-2-SOUND, a training-free method that approaches this specific problem by employing the foundational models, CoDi (Tang et al., 2024), DepthAnything (Yang et al., 2024), and Segment Anything (Kirillov et al., 2023). The pipeline works by explicitly factorizing this generative problem into separate problems: (1) identifying regions of interest in the input; (2) finding the 3D location on the viewing sphere for a visual element; (3) generating audio from the regions of interest, and (4) simulating these multiple mono-audio outputs to generate spatial audio. We then explicitly compose these component models to generate spatial audio conditioned on the visual input (e.g. images, animated images (GIFs), or videos) as shown in Figure 1. As our method is compositional and zero-shot, we hope to see SEE-2-SOUND widely used or adapted for content creation and enhancement as well as accessibility.
@@
By generating spatial audio that accurately corresponds to visual elements, we can significantly enhance the immersion of images and videos, creating more realistic and engaging user experiences with digital content.SEE-2-SOUND presents a step towards true complete video and image generation, by also incorporating non-verbal audio that accompanies visual generation models (Singer et al., 2023a; Brooks et al., 2024; Blattmann et al., 2023a, b; Luo et al., 2023; Wang et al., 2024). Additionally, our pipeline can also be employed on real visual content.
Our contributions can be summarized as follows.
@@ 2. Related works
The audio generation field has been gaining a lot ofmuch attention lately, primarily in models that create a diverse range of of audio content types. From speech synthesis to music generation and sound effects creation, past efforts have explored numerous aspects of auditory information synthesis. Learning algorithms have led to ever-improving techniques in the models for generating high-quality, diverse audio. This surge in interest has fueled research efforts pushing the boundaries of audio generation.
2.1. Conditional Audio Generation
@@
Recent work in conditional audio generation has made significant progress using latent diffusion models (Liu et al., 2023) such as AudioLDM (2) (Liu et al., 2023, 2024) and Make-An-Audio (Huang et al., 2023a). Audiogen (Kreuk et al., 2023) was one of the first to create a text-to-audio environment, treating audio generation as a conditional language modeling task. Building on this, AudioLDM-2 (Liu et al., 2024) introduces a unified audio generation framework that addresses different objectives for speech, music, and sound effects, using the “language of audio” (LoA) and a GPT-2 (Radford et al., 2019) model for translation and latent diffusion (Rombach et al., 2022) for generation.
@@
AudioLM (Borsos et al., 2023) employs neural codecs (Wang et al., 2023) for long-term consistency in audio generation, mapping input audio to discrete tokens and using a language modeling task. Meanwhile, DiffSound (Yang et al., 2023), using a text encoder, decoder, vector quantized variational autoencoder (VQVAE) (Kingma and Welling, 2013), and “vocoder”, introduces a mask-based text generation strategy (MBTG) for generating text descriptions from audio labels to address the scarcity of audio-text paired data. NaturalSpeech (Tan et al., 2024) integrates a memory component in variational autoencoders (Kingma and Welling, 2013) to achieve human-level quality in text-to-speech systems.
Significant progress in text-to-music generation has been made by notable models such as MusicLM (Agostinelli et al., 2023),
@@
Noise2Music (Huang et al., 2023b), MusicGen (Copet et al., 2024) and MeLoDy (Lam et al., 2024). MusicLM (Agostinelli et al., 2023) uses contrastive pretraining to align music and language embeddings, improving performance without explicit text using a semantic modeling stage based on w2v-BERT (Chung et al., 2021). MusicGen (Copet et al., 2024) adopts a language modeling approach with melodic conditioning, enhancing control over output. MeLoDy (Lam et al., 2024) employs dual-path diffusion-based modeling for computational efficiency in music generation.
@@
The image-to-audio and video-to-audio domains have also seen significant advancements. For instance, Im2Wav (Sheffer and Adi, 2023) generates audio from images or image sequences, using Transformers (Vaswani et al., 2017) to manipulate hierarchical discrete audio representation derived from VQ-VAE (Van Den Oord et al., 2017). Similarly, based on GANs (Goodfellow et al., 2020): SpecVQGAN (Iashin and Rahtu, 2021) enhances visually driven audio synthesis from videos, while RVQGAN (Kumar et al., 2024) achieves high-fidelity audio compression. (Aytar et al., 2017) introduces aligned representations that help in cross-modal retrieval and transferring between audio, visual, and text modalities. Furthermore, AV-NeRF (Liang et al., 2024) synthesizes new videos with spatial audio along different camera trajectories, and AV-RIR (Ratnarajah et al., 2023) estimates audio-visual room impulse responses from images and audio.
@@
Multi-modal generative models of note include CoDi(-2) (Blattmann et al., 2023b; Tang et al., 2023). CoDi (Blattmann et al., 2023b) generates multiple output modalities from various input combinations using a composable generation strategy. CoDi-2 (Tang et al., 2023) extends this with enhanced modality-language alignment for encoding and generative tasks, improving zero-shot capabilities in multi-modal generation, reasoning, and interactive conversations.
@@
Recent works like MAGNeT (Ziv et al., 2024) introduce a masked generative sequence modeling method for direct audio token generation, achieving higher efficiency in text-to-music and text-to-audio tasks. DIFF-FOLEY (Luo et al., 2024) improves video-to-audio synthesis quality, relevance, and synchronization using latent diffusion and contrastive audio-visual pretraining. Affusion (Xue et al., 2024) enhances text-to-audio generation by leveraging text-to-image diffusion models, producing high-quality audio closely aligned with text prompts. These advancements significantly expand the possibilities of audio generation.
2.2. Acoustic Matching
@@
Recent advancements in acoustic matching integrate visual and audio cues to enhance realism and immersion in generated content. AViTAR (Chen et al., 2022a) performs visual-acoustic matching using a cross-modal transformer to integrate visual properties into audio and learning from in-the-wild videos. AdVerb (Chowdhury et al., 2023) uses a geometry-aware cross-modal transformer to enhance reverberation by leveraging visual cues, significantly outperforming traditional methods. SoundSpaces 2.0 (Chen et al., 2022b) provides geometry-based audio rendering for 3D environments, supporting tasks like audio-visual navigation and source localization with high fidelity and realism.
2.3. Construction of Surround Sound
@@
In 1881, Ader used a pair of microphones and telephone receivers to relay the sound signals at the Paris Opera (Rumsey, 2012), a first-of-its-kind spatial (stereophonic) sound demonstration. In the past three decades, the development of digital signal processing (Kirkeby et al., 1996), computer, and internet technologies have rapidly evolved spatial sound (Xie, 2020). Together, they have made spatial audio commonplace now (Zhang et al., 2017).
@@
In the context of virtual environments generated by game engines, spatial audio is crucial to provide an immersive experience (Broderick et al., 2018). Traditionally, sounds in virtual worlds are created by recording individual sound sources separately and then blending them according to the intended scene setup (Murphy and Neff, 2011) using various audio processing techniques (Gerzon, 1973; Dickins and Kennedy, 1999). But this recording process is costly (Zhang et al., 2017) and optimizing it involves generating spatial audio. While spatializing mono-audio is a popular area of research (discussed below), to the best of our knowledge, there have not been any approaches to generate spatial audio directly.
2.4. Audio Spatialization
@@
Recent advancements in audio spatialization have focused on integrating audio and visual modalities to enhance realism. Sep-Stereo (Zhou et al., 2020) uses an associative pyramid structure to generate stereo output by combining audio and visual information. Similarly, LAVSS (Ye et al., 2024) introduces a location-guided audio-visual spatial audio separator, highlighting the importance of visual cues in audio generation.
@@
Several studies have synchronized visual cues and encoders to produce natural sound for dynamic video scenarios (Zhou et al., 2018). Methods like those in (Gan et al., 2019; Gao and Grauman, 2019) use a U-Net to encode monaural input and decode binaural output with visual guidance.
@@
Other approaches, such as  (Gan et al., 2020; Kreuk et al., 2023) generate audio samples conditioned on diverse inputs such as text, motion key points, and position information (Garg et al., 2021; Parida et al., 2022), respectively. Techniques like (Morgado et al., 2018) utilize end-to-end trainable neural networks for sound source separation and localization in $360$-degree videos. Additionally, works like (Li et al., 2018) synthesize directional impulse responses by exploiting the spatial and directional characteristics of room reverberation.
Figure 2. Overview of our pipeline. Given an input image we segment objects, estimate depth, generate per-object audio, then we place all of these audio according to the 3D co-ordinates on a viewing sphere producing spatial or surround sound.
[…]
Methods Figure.
2.5. Audio Visual Learning
@@
In recent years, there has been a growing interest in audio-visual learning, focusing on understanding the inherent connection between visual and auditory stimuli within vision research. Noteworthy works have delved into explaining the natural synchronization between visual and audio elements (Arandjelovic and Zisserman, 2017; Aytar et al., 2016; Owens et al., 2016). This approach finds application in diverse domains, including speech recognition utilizing both audio and visual cues (Son Chung et al., 2017; Hu et al., 2016; Yu et al., 2020; Zhou et al., 2019), event localization through joint analysis of audio and visual information (Tian et al., 2018, 2020; Wu et al., 2019), and sound source localization (Arandjelovic and Zisserman, 2018; Hu et al., 2020; Rouditchenko et al., 2019; Senocak et al., 2018; Tian et al., 2018; Zhao et al., 2018). Additionally, there have been explorations in the use of audio-visual data for unsupervised representation learning, as demonstrated in (Owens and Efros, 2018; Owens et al., 2016).
@@
3. Preliminary: Conditional LDM
Latent Diffusion models (LDMs) (Rombach et al., 2022) project input images into a lower-dimensional latent space. Unlike traditional diffusion models that operate in a high-dimensional pixel space, LDMs perform forward and backward diffusion processes in a lower-dimensional learned latent space. For generating audio, we have a conditional LDM ${\bm{\epsilon}}_{\theta}$, which can be interpreted as a sequence of denoising autoencoders ${\bm{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\ldots)$; $t\sim\mathcal{U}({1,\ldots,T})$, trained using a squared error loss (representing the denoising term in the ELBO) to denoise some variably noisy audio signals using the following objective,
(1) $\mathbb{E}_{\mathcal{E}(\mathbf{x}),\mathbf{c},{\bm{\epsilon}}\sim\mathcal{N}(0,I),t}\!\left[|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}(\mathbf{z}_{t},t,\tau_{\theta}(\mathbf{c}))|^{2}_{2}\right],$
@@ 3. Preliminary: Conditional LDM
4. Method
@@ 4.1. Source Estimation
We first focus on finding out regions of interest from the image, $\mathbf{I}$. We do this by finding a segmentation mask ${\mathcal{M}}\in{\mathbb{R}}^{H\times W}$, using a pre-trained Segment Anything model (Kirillov et al., 2023). From the segmentation mask ${\mathcal{M}}$, we extract the contours of each region of interest. Let ${\mathcal{C}}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{n}\}$ represent the set of contours, where each contour $\mathbf{c}_{k}$ is a sequence of points outlining one region of interest. We then compute the smallest polygon ${\mathcal{P}}_{k}$ that bounds each contour $\mathbf{c}_{k}$. Formally, we solve the following optimization problem for each contour,
(2) $\min_{{\mathcal{P}}_{k}}\operatorname{Area}({\mathcal{P}}_{k})\text{ , subjectto, }\forall\mathbf{c}_{k}\in{\mathcal{P}}_{k},$
@@
(3) ${\mathcal{R}}_{k}=\left[\min(\mathbf{c}_{k}^{(x)}),\max(\mathbf{c}_{k}^{(x)}),\min(\mathbf{c}_{k}^{(y)}),\max(\mathbf{c}_{k}^{(y)})\right].$
@@
For the image $\mathbf{I}$, we also estimate the monocular depth map ${\mathcal{D}}\in{\mathbb{R}}^{H\times W}$ using a pre-trained Depth Anything model (Yang et al., 2024). Now we can construct a set of image patches $\mathbf{I}_{\mathbf{p}}$, for the regions of interest from either ${\mathcal{P}}$ or ${\mathcal{R}}$ to create a set of variably sized patches, and the 3D coordinates $(x,y,z)$, from the image coordinates and the monocular depth coordinates illustrated in Figure 2. We find that using AABBs usually works better for generating spatial audio.
4.2. Audio Generation